OpenStudy (anonymous):

What is the simplified form of y^2+5y+6/y^2+2y-3?

OpenStudy (anonymous):

\[y^2+5y+6=(y+3)(y+2)\] \[y^2+2y-3=(y-1)(y+3)\] so that help?

OpenStudy (anonymous):

yes thanks